ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: nonconmmutative geometry, pseudodifferential operators Ponge, Raphael, Orsay (France) Wodzicki Residue for Heisenberg Manifolds We construct a Wodzicki residue for Heisenberg manifolds. Such a manifold M comes with a sub-bundle V of T M such that V and [V, V] span T M. This the case for CR and contact manifolds. We prove that this Wodzicki residue extends the Dixmier trace and is the unique trace on the algebra of pseudodifferential operators we consider.